Records Storage Options

Center City Ambulatory Clinic has upgraded their server system, and it now can currently hold virtually unlimited data. The health information supervisor is considering pulling all of the charts at the storage facility and having staff scan them in to the patient’s charts as “batches” (not individually filing each page, but scanning them in, and if the patient comes back for an appointment, they will file each page to the appropriate location in the chart at that time).

Each chart stored offsite has an average of 40 pages in it. The supervisor believes she can hire one or two part time staff for $10.00 per hour to scan these in. Someone scanning and filing each batch should be able prep, scan and file approximately 30 charts per hour.

  1. How many hours will it take staff to scan in all 40,000 charts? (show your work)
  1. Which option will save this company more money; keeping the records in storage, or hiring someone to scan them in, and why?
Answer #1

Per hour= 30 charts = 30/60

For 40000 = 40000/x

30÷60 = 40000÷ x

30x = 60(40000)

X= 60(40000)÷ 30

X = 80000

To calculate hours= 80000÷ 60= 1333 hours.

1333÷ 24 = 55.5 days

The company can have the document in online option hence all health professional can hire the needed data.If it is in online hire method there is no need to scan and store.Directly can obtain data.